OpenStudy (anonymous):

Find S11 for 1+2+4+8+...

OpenStudy (anonymous):

I have 1(1-2^11)/(1-2) = (1-2048)/-1 = (-2047)/-1 = 2047?

OpenStudy (shubhamsrg):

yes, that is absolutely correct!

OpenStudy (anonymous):

yay!! :) ok thanks